OpenStudy (anonymous):

I'm unsure about my answer for the Molecular and Empirical formula. Can H3PO2 be the empirical formula of H3PO4? And can C3H2 be the empirical formula of C3H8? Or am I not suppose to reduce it if the other subscript can't? If that makes any sense...

OpenStudy (the_fizicx99):

What is this? Physics?

OpenStudy (anonymous):

No its Chemistry.

OpenStudy (the_fizicx99):

Sorry I'm still learning the basics, 9th grader. :(

OpenStudy (anonymous):

Its okay. Thanks for at least looking at it! Its simply reducing so i think i'm over thinking it. Because the other problems were like this: P4O10 to P205

OpenStudy (cwrw238):

the answer is no inj both cases the empirical fomula is the simplest formula which is found by analysis eg the empirical formula of C2H4 is CH2

OpenStudy (the_fizicx99):

Oh I think I have a decent concept, might confuse you so nvm >.> Still ehh with it myself..

OpenStudy (anonymous):

I know I'm over thinking it but cwrw238 the example you gave has two even subscripts. Mine don't. So does that matter or not? Does it go from C3H8 to C3H2? Is that right or do i leave it alone?

OpenStudy (anonymous):

isnt there mols involved ?

OpenStudy (anonymous):

Umm idk haha. My teacher gave us the molecular formula and we are suppose to write the empirical formula. Its like a chart

OpenStudy (anonymous):

do you know the mass or the percent of each element ?

OpenStudy (anonymous):

I could figure it out but will that help me find the empirical formula?

OpenStudy (anonymous):

yes , im looking over my notes so i can help but what will you figure out ? the percent ?

OpenStudy (anonymous):

I'll do anything to help you help me figure it out xD

OpenStudy (anonymous):

haha ok ill try my hardest to explain , ik how to do it i just dont know what every step is called alright ?

OpenStudy (anonymous):

okay lets say you know the percent of each element , ima do an example ok ?

OpenStudy (nikato):

Omg. I'm doing this too! But shouldn't u know some sort of weight of like the entire formula in order to find the empirical formula

OpenStudy (anonymous):

Okay thats what i thought... I was getting confused but your sweet wednylisette!

OpenStudy (anonymous):

92.26% carbon , 7.74 % hydrogen carbons atomic mass = 12.01 hydrogens atomic mass = 1 so your going to divide the percent by its mass carbon = \[\frac{ 92.26 }{ 12.01 }\] hydrogen = \[\frac{ 7.74 }{ 1 }\]

OpenStudy (anonymous):

you get it so far ?

OpenStudy (anonymous):

Yeah but i don't get how it goes back to the Molecular and Empirical formula question i asked haha :)

OpenStudy (anonymous):

just watch lol

OpenStudy (anonymous):

okay after you divide them carbon = 7.681 and hydrogen = 7.74 this next part will tell you what subscripts your empirical formula will have

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

your going to take the smallest number . 7.681 is smaller than 7.74 so you will divide BOTH numbers by the smallest number which is 7.681 .

OpenStudy (anonymous):

carbon 7.681 / 7.681 = 1 hydrogen 7.74 / 7.681 = 1.007 which is basically 1

OpenStudy (anonymous):

so carbon is 1 and hydrogen is 1 your empirical formula is now C1H1 = CH

OpenStudy (anonymous):

And thats for C3H8? Just making sure we're on the same page.

OpenStudy (anonymous):

or that was just your example?

OpenStudy (anonymous):

no thats an example . i can only figure it out if you give me the mols or the percent

OpenStudy (anonymous):

but ok lets do C3H8 lets find the percent

OpenStudy (anonymous):

alright. I'll do one and you do one. I'll do C3 and you do H8, is that okay?

OpenStudy (anonymous):

you know how to get the percents ?

OpenStudy (anonymous):

I'm looking haha. I did some earlier today but I don't think its the same way you did.

OpenStudy (anonymous):

okay well find the atomic mass of each C3 means theres 3 carbons so you will multiply the atomic mass by three , theres 8 hydrogens so you will multiply the atomic mass of hydrogen by 8.

OpenStudy (anonymous):

C 12.01(3) = 36.03 H 1 ( 8 ) = 8

OpenStudy (anonymous):

okay I got that far it was after that i didn't know.

OpenStudy (anonymous):

now you just gotta find the percent so add those two numbers up 36.03 + 8 = 44.03

OpenStudy (anonymous):

and then you divide them both by 8?

OpenStudy (anonymous):

now your gonna take the molar mass ( which is what we did in step one ) and divide it by the formula mass which is 44.03 ( the sum of both masses )

OpenStudy (anonymous):

example with carbon . 36.03 / 44.03 = 0.8183 to turn that into a percent multiply it by 100 so now carbon is 81.83 %

OpenStudy (anonymous):

Hydrogen 8/44.03 = 0.18169 x 100 = 18.169 which you can round to 18.17 %

OpenStudy (anonymous):

ok... so I think I got it now. I think the answer is just stays the same. So its C3H8

OpenStudy (anonymous):

now you gotta do the other part though , like what i did with my example

OpenStudy (anonymous):

Okay thanks! I think i got it now :)

OpenStudy (anonymous):

i hope so (:

OpenStudy (the_fizicx99):

Nice explanation @wendylisette what grade are you both in?

OpenStudy (anonymous):

i'm in 10th although i'm the age to be in 11th... bad homeschooling got me behind XD

OpenStudy (anonymous):

And yet years later i'm homeschooling again

OpenStudy (anonymous):

10th